OpenStudy (zepp):

Doing some calculus questions... And, what's the difference between 'Differentiate each function' and 'Determine each derivative'?

OpenStudy (zepp):

1. Differentiate each function. a) \(y=x^5\) b) \(y=x^{10}\) ..and so on 2. Determine each derivative a) \(\large\frac{d}{dx}(x^9)\) b) \(\large \frac{d}{dx}(x^{20})\) ...and so on Aren't they the same? D:

OpenStudy (zepp):

@experimentX Help please :)

OpenStudy (experimentx):

same type .. follow this rule. this is called power rule. \[ \large \frac{d}{dx}(x^{n}) = n \times x^{n-1}\]

OpenStudy (zepp):

Alright, thanks :) I know the power rule, but I just want to know the difference between the two questions. (So there's none, lol)

OpenStudy (experimentx):

well ... the questions are different as the function are different but the type of question is same.

OpenStudy (zepp):

Okay cool, thanks! :)

OpenStudy (experimentx):

yw